### v1

# Adaptive Incentive Design with Learning Agents

Abstract
How can the system operator learn an incentive mechanism that achieves social optimality based on limited information about the agents’ behavior, who are dynamically updating their strategies? To answer this question, we propose an adaptive incentive mechanism. This mechanism updates the incentives of agents based on the feedback of each agent’s externality, evaluated as the difference between the player’s marginal cost and society’s marginal cost at each time step. The proposed mechanism updates the incentives on a slower timescale compared to the agents’ learning dynamics, resulting in a two-timescale coupled dynamical system. Notably, this mechanism is agnostic to the specific learning dynamics used by agents to update their strategies. We show that any fixed point of this adaptive incentive mechanism corresponds to the optimal incentive mechanism, ensuring that the Nash equilibrium coincides with the socially optimal strategy. Additionally, we provide sufficient conditions that guarantee the convergence of the adaptive incentive mechanism to a fixed point. Our results apply to both atomic and non-atomic games. To demonstrate the effectiveness of our proposed mechanism, we verify the convergence conditions in two practically relevant games: atomic networked quadratic aggregative games and non-atomic network routing games.
1 Introduction
Incentive mechanisms play a crucial role in many societal systems, where the outcomes are governed by the selfish interactions of a large number users (or algorithms on their behalf) and the system operator. The outcome of such strategic interactions, characterized by the Nash equilibrium, is often sub-optimal because of the fact that individual agents often do not account for the externality of their actions (i.e., how their actions affect the costs of others), when minimizing their own costs. An important way to address the sub-optimality is to provide agents with incentives that align their individual goal of cost minimization with the goal of minimizing the total cost of the societal system[Lev85, HLO82, Baş84].
The system operator faces two key challenges when designing such incentive mechanisms. First, due to the system’s scale and privacy concerns, the designer often has limited information about individual agents’ costs and the cost function of the entire system. Second, the system operator often needs to consider the learning behavior of agents, who repeatedly update their strategies in response to the incentive mechanism, especially when the physical system experiences a random shock and players are trying to reach a new equilibrium [BG14, CM21, MKWS21]. As a result, the designer cannot infer the equilibrium to predict agents’ behavior, making incentive design a challenging task. This leads to the following question:
How can the system operator learn an incentive mechanism that achieves social optimality based on limited information about the behavior of agents who are adaptively updating their strategies?
To answer this question, we propose an adaptive incentive mechanism that updates the incentives based on the behavior of agents, which is repeatedly updated as part of some learning process, and the gradient information of the cost functions. This results in a coupled dynamical system comprising incentive and strategy updates. Our proposed framework can be applied to both atomic and non-atomic games.
Our proposed incentive mechanism has three key features. First, the incentive update incorporates the externality incurred by the agents’ current strategy, which is evaluated as the difference between their own marginal cost and the marginal cost for the entire system. This feature guarantees that any fixed point of the coupled learning dynamics induces a socially optimal outcome: the incentive provided to each agent equals their externality given their equilibrium strategy, and the associated Nash equilibrium is a socially optimal strategy (Proposition 3.1). Second, the incentive mechanism only requires limited information about the behavior of agents. Specifically, it is agnostic to the strategy update dynamics used by agents and only requires information about the gradient of the cost function of agents at the current strategy to evaluate the externality (Remark 2.2). Third, the incentive update occurs at a slower timescale compared to the agents’ strategy updates. This slower evolution of incentives is a desirable characteristic because frequent incentive updates may hamper agents’ participation. It allows agents to consider the incentives as static while updating their strategies.
We provide sufficient conditions on the underlying game such that the set of fixed points is a singleton, making the socially optimal incentive mechanism unique (Proposition 3.1). We also characterize sufficient conditions that guarantee the convergence of the coupled dynamical system (Proposition 3.3). Since the convergent strategy profile and incentive mechanism correspond to a fixed point that is also socially optimal, these sufficient conditions ensure that the dynamic incentive mechanism eventually induces a socially optimal outcome in the long run. To prove this result, we leverage the timescale separation between the strategy update and the incentive update, using results from the theory of two-timescale dynamical systems [Bor97]. This theory allows us to analyze the convergence of the coupled dynamical system in two steps by decoupling the convergence of the strategy update and the incentive mechanism update. First, the convergence of the strategy update (on the faster timescale) is analyzed by viewing the incentive mechanism (on the slower timescale) as static. In our work, we off-load this convergence to the rich literature on learning in games (e.g., [FL98, San10, MS96]). Second, the convergence of the incentive mechanism update is analyzed via the associated continuous-time dynamical system, evaluated at the converged value of the strategy update (i.e., Nash equilibrium). The sufficient conditions developed in Proposition 3.3 are built on results on the convergence of non-linear dynamical systems from cooperative dynamical systems theory [Hir85] and Lyapunov-based methods [Sas13].
To demonstrate the usefulness of the adaptive incentive mechanism, we apply it to two classes of practically relevant games: (i) atomic networked quadratic games and (ii) non-atomic routing games. In atomic networked quadratic games, each agent’s cost function depends on their own strategy as well as the aggregate strategies of their opponents. This aggregation is done through a linear combination of the neighboring agents’ strategies with weights characterized by a network matrix. Our proposed incentive mechanism enables the system operator to adaptively update incentives, based on the externality of each agent’s strategy on their neighbors, while the agents are learning the equilibrium strategies. Our results when specialized to this setup provide sufficient conditions on the network matrix to ensure that the system converges to a socially optimal outcome.
Furthermore, in the non-atomic routing games, agents (travelers) make routing decisions in a congested network with multiple origin-destination pairs. The system operator imposes incentives in the form of toll prices on each network edge. Our proposed incentive mechanism adaptively gets updated using only the observed flow on edges and the gradient of the edge latency function. Agents can adopt a variety of strategy updates that lead to the equilibrium of the routing game. We show that the adaptive incentive mechanism converges to the toll price that minimizes total congestion.
The article is organized as follows: in Sec. 2, we describe the setup of both atomic and non-atomic game considered here. In addition, we also provide the joint strategy and incentive update considered in this paper. We present the main results on the fixed point and convergence of adaptive incentive mechanism in Sec. 3. In Sec. 4, we study the application of our proposed adaptive incentive mechanism on the atomic networked aggregative games (Sec. 4.1) and non-atomic routing game (Sec. 4.2). We conclude our work in Sec. 5.
1.1 Related Works
Two-timescale Learning Dynamics: Learning dynamics where incentives are updated at a slower timescale than the strategy of agents have been studied in [MNPQ22, COMP23, PBMT17, OP22, RF20, LLY+21, LNL+23]. Specifically, [MNPQ22] studies Stackelberg games with a single leader and a population of followers. In their approach the leader deploys gradient-based update and followers update their strategies using replicator dynamics. [COMP23] and [PBMT17] consider incentive design in affine congestion games, where the incentives are updated using a distributed version of gradient descent. [OP22] studies incentive design to control traffic congestion on a single highway using gradient based incentive updates. [RF20] studies the problem of incentive design while learning the cost functions of agents. The authors assume that both the cost functions and incentive policies are linearly parameterized, and the incentive updates rely on the knowledge of agents’ strategy update rules instead of just the current strategy as in our setting. Additionally, [LLY+21, LNL+23] propose a two-timescale discrete-time learning dynamics, where agents adopt a mirror descent-based strategy update, and the social planner updates the incentive parameter according to a gradient based method. The convergence of these gradient-based learning dynamics relies on the assumption that the social cost given agents’ equilibrium strategy is convex in the incentive parameter. However, the convexity assumption is restrictive as highlighted by an example in Appendix A.
Same-timescale Learning Dynamics: The problem of steering non-cooperative agents to a desired Nash equilibrium by considering an incentive update which runs on the same timescale as strategy updates is studied in [SCM22a, SCM22b, ZFA+23]. Particularly, in [SCM22a] the authors study such updates in a specific setting of quadratic aggregative games. In [SCM22b], the authors consider a setup where the agents’ cost is only dependent on their own action and the price signal which is provided by operator which is used by agents based on their individual reliability over the operator. The aim is to design this price signal so that the reliability on the price signal is retained while steering the equilibrium to the desired point. In [ZFA+23], the authors consider the problem of steering no-regret learning agents to optimal equilibrium but the computation of incentive mechanism requires solving an optimization problem at each time step.
Learning in Stackelberg Games: Our work is also related to the literature on learning in Stackelberg games, where the planner often has limited information about the game between agents and need to design optimal mechanism by dynamically receiving feedback from agents’ responses (for instance, refer to [BHP14, LCM09, PSTZ19, BJWX21, CFD24, JGA+22, MSRM23]). Typically, this line of research posit structural assumptions on the game between followers (e.g., finite action space of agents, or linearly parametrized utility function of agents)[BHP14, LCM09, PSTZ19, BJWX21, CFD24, JGA+22], or only ensures convergence to a locally optimal solution [MSRM23].
In comparison to the preceding three lines of research, our approach of externality-based incentive design is applicable for a broad class of games (with continuous action spaces and potentially non-linear utility functions), while providing sufficient conditions to ensure asymptotic convergence to the optimal solution.
Notations
Given a function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, we use $\nabla_{x_{i}}f(x)$ to denote $\frac{\partial f}{\partial x_{i}}(x)$, the derivative of $f$ with respect to $x_{i}$ for any $i\in\{1,2,...,n\}$ and $\nabla f$ to be the Jacobian of the function. For any set $A$ we use $\textsf{conv}(A)$ to denote the convex hull of the set. For any set $X\subseteq\mathbb{R}^{n}$, we call a function $f:X\rightarrow\mathbb{R}$ to be Lipschitz if there exists a positive scalar $L$ such that $\|f(x)-f(x^{\prime})\|\leqslant L\|x-x^{\prime}\|$ for every $x,x^{\prime}\in X$. For any vector $x\in X$ and any positive scalar $r>0$, the set $\mathcal{B}_{r}(x)=\{x^{\prime}\in X|\|x^{\prime}-x\|<r\}$ denotes the $r$-radius neighborhood of the vector $x$. For any set $X$, we define $\textsf{boundary}(X)$ and $\textsf{int}(X)$ to be the boundary and interior of set $X$, respectively. Finally, for any function $f(\cdot)$, we denote the domain of the function by $\textsf{dom}(f)$. For any vector $x\in\mathbb{R}^{n}$, we define $\textsf{diag}(x)\in\mathbb{R}^{n\times n}$ to be a diagonal matrix with diagonal entries corresponding to $x$.
2 Model
We introduce both atomic and non-atomic static games in Sec. 2.1. In Sec. 2.2, we present the adaptive incentive design approach proposed in this work.
2.1 Static Games
2.1.1 Atomic Games
Consider a game $G$ with a finite set of players $\mathcal{I}$. The strategy of each player $i\in\mathcal{I}$ is ${x}_{i}\in X_{i}$, where $X_{i}$ is a non-empty and closed interval in $\mathbb{R}$. The joint strategy profile of all players is ${x}=({x}_{i})_{i\in\mathcal{I}}$, and the set of all joint strategy profiles is $X\coloneqq\prod_{i\in\mathcal{I}}X_{i}$. The cost function of each player $i\in\mathcal{I}$ is $\ell_{i}:\mathbb{R}^{|\mathcal{I}|}\to\mathbb{R}$ [^1] We measure the outcome of our games by costs instead of utilities. Equivalently, the utility of each player is the negative value of the cost.
A social planner designs incentives by setting a payment $p_{i}x_{i}\in\mathbb{R}$ for each player $i$ that is linear in their strategy $x_{i}$. Here, $p_{i}\in\mathbb{R}$ represents the marginal payment for every unit increase in strategy of player $i$. The value of $p_{i}x_{i}$ can either be negative or positive, which represents a marginal subsidy or a marginal tax, respectively. Given the incentive mechanism $p=(p_{i})_{i\in I}$, the total cost of each player $i\in\mathcal{I}$ is:
$c_{i}({x},p)=\ell_{i}({x})+p_{i}{x}_{i},\quad\forall\ {x}\in X.$ (1)
A strategy profile $x^{\ast}(p)\in X$ is a Nash equilibrium in the atomic game $G$ with the incentive mechanism $p$ if
$c_{i}(x^{\ast}_{i}(p),x^{\ast}_{-i}(p),p)\leqslant c_{i}({x}_{i},x^{\ast}_{-i}(p),p),~{}\forall\ {x}_{i}\in X_{i},~{}\forall i\in\mathcal{I}.$
A strategy profile ${x}^{\dagger}\in X$ is socially optimal if it minimizes the social cost function $\Phi:\mathbb{R}^{|\mathcal{I}|}\to\mathbb{R}$ over $X$.
Assumption 2.1.
The cost function $\ell_{i}({x}_{i},{x}_{-i})$ is strongly convex in ${x}_{i}$ for all $i\in\mathcal{I}$ and all ${x}_{-i}=({x}_{j})_{j\in\mathcal{I}\setminus\{i\}}$. Moreover, $\ell_{i}(\cdot)$ is continuously differentiable and its gradient is Lipschitz. Additionally, the social cost function $\Phi(x)$ is twice continuously differentiable, has Lipschitz gradient, and is strictly convex.
Under Assumption 2.1, it is known that, for any $p\in\mathbb{R}^{|\mathcal{I}|}$, the Nash equilibrium $x^{\ast}(p)$ exists, is unique and is Lipscthiz continuous in $p$ [Daf88]. Moreover, the socially optimal strategy $x^{\dagger}$ is also unique.
2.1.2 Non-atomic Games
Consider a game $\tilde{G}$ with a finite set of player populations $\tilde{\mathcal{I}}$. Each population $i\in\tilde{\mathcal{I}}$ is comprised of a continuum set of (infinitesimal) players with mass $\tilde{M}_{i}>0$. Every (infinitesimal) player in any population can choose an action in a finite set $\tilde{S}_{i}$. The strategy distribution of population $i\in\tilde{\mathcal{I}}$ is $\tilde{x}_{i}=(\tilde{x}_{i}^{j})_{j\in\tilde{S}_{i}}$, where $\tilde{x}_{i}^{j}$ is the mass of players in population $i$ who choose action $j\in\tilde{S}_{i}$. Then, the set of all strategy distributions of population $i$ is $\tilde{X}_{i}=\left\{\tilde{x}_{i}\in\mathbb{R}^{|\tilde{S}_{i}|}|\sum_{j\in\tilde{S}_{i}}\tilde{x}_{i}^{j}=\tilde{M}_{i},~{}\tilde{x}_{i}^{j}\geqslant 0,\forall j\in\tilde{S}_{i}\right\}$. The strategy distribution of all populations is $\tilde{x}=(\tilde{x}_{i})_{i\in\tilde{\mathcal{I}}}\in\tilde{X}=\prod_{i\in\tilde{\mathcal{I}}}\tilde{X}_{i}$. We define $\tilde{S}=\prod_{i\in\tilde{\mathcal{I}}}\tilde{S}_{i}$. Given a strategy distribution $\tilde{x}\in\tilde{X}$, the cost of players in population $i\in\tilde{\mathcal{I}}$ for choosing action $j\in\tilde{S}_{i}$ is ~$\tilde{\ell}_{i}^{j}(\tilde{x})$. We denote ~$\tilde{\ell}_{i}(\tilde{x})=(\tilde{\ell}_{i}^{j}(\tilde{x}))_{j\in\tilde{S}_{i}}$ as the vector of costs for each $i\in\tilde{\mathcal{I}}$.
A social planner designs incentives by setting a payment $\tilde{p}_{i}^{j}$ for players in population $i$ who choose action $j\in\tilde{S}_{i}$. Consequently, given the incentive mechanism $\tilde{p}=({\tilde{p}_{i}^{j}})_{j\in S_{i},i\in\tilde{\mathcal{I}}}$, the total cost experience by any player, in population $i\in\tilde{\mathcal{I}}$, that chooses action $j\in\tilde{S}_{i}$ is
$\tilde{c}_{i}^{j}(\tilde{x},\tilde{p})=\tilde{\ell}_{i}^{j}(\tilde{x})+\tilde{p}_{i}^{j},\quad\forall\ \tilde{x}\in\tilde{X}.$ (2)
A strategy distribution ${\tilde{x}^{\ast}}(\tilde{p})\in\tilde{X}$ is a Nash equilibrium in the nonatomic game $\tilde{G}$ with $\tilde{p}$ if
${\tilde{x}^{j\ast}_{i}}(\tilde{p})$ $>0,\quad$
$\Rightarrow\tilde{c}_{i}^{j}({\tilde{x}^{\ast}}(\tilde{p}),\tilde{p})\leqslant\tilde{c}_{i}^{j^{\prime}}({\tilde{x}^{\ast}}(\tilde{p}),\tilde{p}),~{}\forall j,j^{\prime}\in\tilde{S}_{i},~{}\forall i\in\tilde{\mathcal{I}}.$
A strategy distribution ${\tilde{x}}^{\dagger}\in\tilde{X}$ is socially optimal if ${\tilde{x}}^{\dagger}$ minimizes a social cost function $\tilde{\Phi}:\mathbb{R}^{|\tilde{S}|}\to\mathbb{R}$.
Assumption 2.2.
The function $\tilde{\ell}(\cdot)$ is Lipschitz continuous. Moreover, it is strongly monotone: there exists $\rho>0$ such that
~$\left\langle\tilde{\ell}(\tilde{x})-\tilde{\ell}(\tilde{x}^{\prime}),\tilde{x}-\tilde{x}^{\prime}\right\rangle>\rho\|\tilde{x}-\tilde{x}^{\prime}\|^{2},\quad\forall\tilde{x}\neq\tilde{x}^{\prime}\in\tilde{X}.$
Additionally, $\tilde{\Phi}(\tilde{x})$ is twice continuously differentiable, has Lipschitz gradient, and strictly convex.
Under Assumption 2.2, Nash equilibrium $x^{\ast}(p)$ exists, is unique and Lipscthiz continuous for every $p\in\mathbb{R}^{|\tilde{\mathcal{I}}|}$ [San10, Daf88]. Moreover, the socially optimal strategy distribution is also unique.
2.2 Coupled Strategy and Incentive Update
We consider a coupled dynamics that jointly update agents’ strategies and the incentive mechanism with discrete timesteps $k\in\mathbb{N}$. At step $k$, the strategy profile in the atomic game $G$ (resp. non-atomic game $\tilde{G}$) is $x_{k}=(x_{i,k})_{i\in\mathcal{I}}$ (resp. $\tilde{x}_{k}=(\tilde{x}_{i,k})_{i\in\tilde{\mathcal{I}}}$), where $x_{i,k}$ (resp. $\tilde{x}_{i,k}$) is the strategy of player $i$ (population $i$), and the incentive mechanism is $p_{k}=(p_{i,k})_{i\in\mathcal{I}}$ (resp. $\tilde{p}_{k}=(\tilde{p}_{i,k}^{j})_{j\in S_{i},i\in\tilde{\mathcal{I}}}$). The strategy updates and the incentive updates are presented below:
Strategy update.
$x_{k+1}$ $=(1-\gamma_{k})x_{k}+\gamma_{k}f(x_{k},p_{k}),$ ($x$-update)
$\tilde{x}_{k+1}$ $=(1-\gamma_{k})\tilde{x}_{k}+\gamma_{k}\tilde{f}(\tilde{x}_{k},\tilde{p}_{k}).$ ($\tilde{x}$-update)
In each step $k+1$, the updated strategy is a linear combination of the strategy in stage $k$ (i.e. $x_{k}$ in $G$ and $\tilde{x}_{k}$ in $\tilde{G}$), and a new strategy $f(x_{k},p_{k})\in X$ in $G$ (resp. $\tilde{f}(\tilde{x}_{k},\tilde{p}_{k})\in\tilde{X}$ in $\tilde{G}$) that depends on the previous strategy and the incentive mechanism. The relative weight in the linear combination is determined by the step-size $\gamma_{k}\in(0,1)$. We require that for any $p$ (resp. $\tilde{p}$), the fixed point associated with update ($x$-update) (resp. ($\tilde{x}$-update)) is a Nash equilibrium, i.e.
$x^{\ast}(p)$ =$=\{x:f(x,p)=x\},\quad\forall p\in\mathbb{R}^{|\mathcal{I}|},$ (3)
$\tilde{x}^{\ast}(\tilde{p})$ =$=\{\tilde{x}:\tilde{f}(\tilde{x},\tilde{p})=\tilde{x}\},\quad\forall\ \tilde{p}\in\mathbb{R}^{|\tilde{\mathcal{I}}|}.$
Some of the examples of commonly studied learning dynamics ($x$-update) and ($\tilde{x}$-update) are delineated below:
  1. 1.
    Equilibrium update ([CFD24, JGA+22]): The strategy update incorporates a Nash equilibrium strategy profile with respect to the incentive mechanism in step $k$:
    $f({x}_{k},p_{k})=x^{\ast}(p_{k}),\quad\text{and}~{}\tilde{f}(\tilde{x}_{k},\tilde{p}_{k})={\tilde{x}^{\ast}}(\tilde{p}_{k}).$ (4)
  2. 2.
    Best response update ([FL98, HS02]): The strategy update incorporates a best response strategy with respect to the strategy and the incentive mechanism in step $k$:
    $f_{i}(x_{k},p_{k})$ $=\underset{y_{i}\in X_{i}}{\arg\min}\ c_{i}(y_{i},x_{-i,k},p_{k}),$ (5)
    $\tilde{f}_{i}(\tilde{x}_{k},\tilde{p}_{k})$ $=\underset{\tilde{y}_{i}\in\tilde{X}_{i}}{\arg\min}\ \tilde{y}_{i}^{\top}\tilde{c}_{i}(\tilde{x}_{k},\tilde{p}_{k}).$
  3. 3.
    Gradient-based update ([LW94, LM13, San10]): The strategy update incorporates a smoothed strategy update with respect to the strategy and the incentive mechanism in step $k$:
    $f_{i}(x_{k},p_{k})$ $=\underset{y_{i}\in X_{i}}{\arg\min}\ z_{i}(x_{k},p_{k})y_{i}-h(y_{i}),$ (6)
    $\tilde{f}_{i}(\tilde{x}_{k},\tilde{p}_{k})$ $=\underset{\tilde{y}_{i}\in\tilde{X}_{i}}{\arg\min}\ \tilde{c}_{i}(\tilde{x}_{k},\tilde{p}_{k})-\tilde{h}(\tilde{y}_{i}),$
    where $z_{i}(x_{k},p_{k})=x_{k}-\eta\nabla_{x_{i}}c_{i}(x_{k},p_{k})$, $\eta$ is step size, and $h(\cdot),\tilde{h}(\cdot)$ are regularizers. If $h(\cdot)$ is a quadratic function then the update becomes projected gradient descent [MZ19]. Furthermore, if $\tilde{h}(\cdot)$ is the entropy function then the update becomes a perturbed best-response update[San10].
Incentive update.
$p_{k+1}$ $=(1-\beta_{k})p_{k}+\beta_{k}e(x_{k}),$ ($p$-update)
$\tilde{p}_{k+1}$ $=(1-\beta_{k})\tilde{p}_{k}+\beta_{k}\tilde{e}(\tilde{x}_{k}),$ ($\tilde{p}$-update)
where $e(x)=(e_{i}(x))_{i\in\mathcal{I}}$, $\tilde{e}(\tilde{x})=(\tilde{e}_{i}^{j}(\tilde{x}))_{j\in S_{i},i\in\tilde{\mathcal{I}}}$, and
$e_{i}({x})$ $=\nabla_{{x}_{i}}\Phi({x})-\nabla_{{x}_{i}}\ell_{i}({x}),\quad\forall i\in\mathcal{I},$ (7a)
$\tilde{e}_{i}^{j}(\tilde{x})$ $=\nabla_{\tilde{x}_{i}^{j}}\tilde{\Phi}(\tilde{x})-\tilde{\ell}_{i}^{j}(\tilde{x}),\quad\forall j\in S_{i},\quad\forall i\in\tilde{\mathcal{I}}.$ (7b)
In (7a), $e_{i}({x})$ is the difference between the marginal social cost, and the marginal cost of player $i$ given ${x}$. Similarly, $\tilde{e}_{i}^{j}(\tilde{x})$ is the difference between the marginal social cost, and the cost experienced by the players in population $i$ who chooses action $j$. We refer $e_{i}({x})$ and $\tilde{e}_{i}(\tilde{x})=(\tilde{e}_{i}^{j}(\tilde{x}))_{j\in S_{i}}$ as the externality of player $i$ and population $i$, respectively, since they capture the difference in impacts of their strategy on the social cost and individual cost.
The updates ($p$-update)-($\tilde{p}$-update) modify the incentives on the basis of the externality caused by the players. In each step $k+1$, the updated incentive mechanism is a linear combination of the incentive mechanism in step $k$ (i.e. $p_{k}$ in $G$ and $\tilde{p}_{k}$ in $\tilde{G}$), and the externality (i.e. $e(x_{k})$ in $G$ and $\tilde{e}(\tilde{x}_{k})$ in $\tilde{G}$) given the strategy in step $k$. The relative weight in the linear combination is determined by the step size $\beta_{k}\in(0,1)$.
In summary, the joint evolution of strategy and incentive mechanism $(x_{k},p_{k})_{k=1}^{\infty}$ (resp. $(\tilde{x}_{k},\tilde{p}_{k})_{k=1}^{\infty}$) in the atomic game $G$ (resp. non-atomic game $\tilde{G}$) is governed by the learning dynamics ($x$-update)–($p$-update) (resp. ($\tilde{x}$-update)–($\tilde{p}$-update)). The step-sizes $(\gamma_{k})_{k=1}^{\infty}$ and $(\beta_{k})_{k=1}^{\infty}$ determine the speed of strategy updates and incentive updates.
Remark 2.1.
Our externality-based incentive update contrasts with the gradient-based incentive update considered in the literature [LLY+21]. Such a gradient-based incentive update typically requires that the equilibrium social cost function $\Phi(x^{\ast}(p))$ (resp. $\tilde{\Phi}(\tilde{x}^{\ast}(\tilde{p}))$) be strongly convex in the incentive mechanism $p$ (resp. $\tilde{p}$). However, this requirement is too restrictive and is not satisfied even in simple games due to the fact that the equilibrium strategy vector is typically non-convex in the incentive mechanism. In Appendix A, we provide an example of a two-link routing game, where the equilibrium social cost function is non-convex in the incentive mechanism, and the gradient-based learning dynamics do not converge.
Remark 2.2.
The incentive update is based on only limited information about the players’ behavior. Specifically, it relies solely on the gradients of the social cost and those of the players’, evaluated at the players’ current strategy profile. Note that it is agnostic to the strategy update dynamics (i.e., ($x$-update), and ($\tilde{x}$-update)) employed by players.
3 General results
In Sec. 3.1, we characterize the set of fixed points of the updates ($x$-update)-($p$-update) and ($\tilde{x}$-update)-($\tilde{p}$-update), and show that any fixed point corresponds to a socially optimal incentive mechanism such that the induced Nash equilibrium strategy profile minimizes the social cost. In Sec. 3.2, we provide a set of sufficient conditions that guarantee the convergence of strategy and incentive updates. Under these conditions, our adaptive incentive mechanism eventually induces a socially optimal outcome.
3.1 Fixed point analysis
We first characterize the set of fixed points of the updates ($x$-update)-($p$-update), and ($\tilde{x}$-update)-($\tilde{p}$-update) as follows:
Atomic game $\text{Atomic game $G$,}~{}\left\{(x,p)|f(x,p)=x,~{}e(x)=p\right\},$ (8a)
[…]
Using (3), from (8a) – (8b), we can write the set of incentive mechanisms at the fixed point, $P^{\dagger}$ as follows:
Atomic game $\text{Atomic game $G$,}~{}P^{\dagger}=\{(p^{\dagger}_{i})_{i\in\mathcal{I}}|e(x^{\ast}(p^{\dagger}))=p^{\dagger}\},$ (9)
[…]
That is, at any fixed point, the incentive of each player is set to be equal to the externality evaluated at their equilibrium strategy profile.
Our first result characterizes conditions under which the fixed point set $P^{\dagger}$ (resp. $\tilde{{P}}^{\dagger}$) is non-empty and singleton in $G$ (resp. $\tilde{G}$). Moreover, given any fixed point incentive parameter $p^{\dagger}\in P^{\dagger}$ and $\tilde{p}^{\dagger}\in\tilde{{P}}^{\dagger}$, the corresponding Nash equilibrium is socially optimal.
Proposition 3.1.
Let Assumptions 2.1 hold and the strategy set $X$ in an atomic game $G$ be compact. The set $P^{\dagger}$ is a non-empty singleton set. The unique $p^{\dagger}\in P^{\dagger}$ is socially optimal, i.e. $x^{\ast}(p^{\dagger})={x}^{\dagger}$.
Moreover, in a non-atomic game $\tilde{G}$ under Assumptions 2.2, $\tilde{{P}}^{\dagger}$ is a non-empty singleton set. The unique $\tilde{p}^{\dagger}\in\tilde{{P}}^{\dagger}$ is socially optimal, i.e., ${\tilde{x}^{\ast}}(\tilde{p}^{\dagger})={\tilde{x}}^{\dagger}$.
Proof.
First, we show that $P^{\dagger}$ is non-empty. That is, there exists $p^{\dagger}$ such that $e(x^{\ast}(p^{\dagger}))=p^{\dagger}$. Define a function $\theta(p)=e(x^{\ast}(p))$. We note that Assumption 2.1 guarantees that $\theta$ is well defined. Thus the problem is reduced to showing existence of fixed point of $\theta(\cdot)$, for which we use the Schauder fixed point theorem [Sma80].
We note that Assumption 2.1 ensures that $\theta(p)$ is a continuous function. Now, define $K\coloneqq\{\theta(p):p\in\mathbb{R}^{|\mathcal{I}|}\}\subseteq\mathbb{R}^{|\mathcal{I}|}$. We claim that the set $K$ is compact. Indeed, this follows by two observations. First, the externality function $e(\cdot)$ is continuous. Second, the range of the function $x^{\ast}(\cdot)$ is $X$ which is a compact set. These two observations ensure that $\theta(p)=e(x^{\ast}(p))$ is a bounded function. Let $\tilde{K}\coloneqq\textsf{conv}(K)$ be the convex hull of $K$, which in turn is also a compact set. Let’s denote the restriction of function $\theta$ on the set $\tilde{K}$ as $\theta_{|\tilde{K}}:\tilde{K}\longrightarrow\tilde{K}$ where $\theta_{|\tilde{K}}(p)=\theta(p)$ for all $p\in\tilde{K}$. We note that $\theta_{|\tilde{K}}$ is a continuous function from a convex compact set to itself and therefore, the Schauder fixed point theorem ensures that there exists $p^{\dagger}\in\tilde{K}$ such that $p^{\dagger}=\theta_{|\tilde{K}}(p^{\dagger})=\theta(p^{\dagger})$. This concludes the proof of existence of $p^{\dagger}$. A similar line of arguments as above can be replicated for non-atomic game $\tilde{G}$ to show that $\tilde{P}^{\dagger}$ is non-empty. We omit the details here for the sake of brevity.
Next, we show that the incentive $p^{\dagger}$ aligns the Nash equilibrium with social optimality (i.e. for any $p^{\dagger}\in P^{\dagger}$, $x^{\ast}(p^{\dagger})={x}^{\dagger}$). Fix $p^{\dagger}\in P^{\dagger}$. For every $i\in\mathcal{I}$ we have $p^{\dagger}_{i}=e_{i}(x^{\ast}(p^{\dagger}))$. This implies $\nabla_{{x}_{i}}\ell_{i}(x^{\ast}(p^{\dagger}))+p^{\dagger}_{i}=\nabla_{{x}_{i}}\Phi(x^{\ast}(p^{\dagger}))$ for every $i\in\mathcal{I}$. This implies
$J(x^{\ast}(p^{\dagger}),p^{\dagger})=\nabla\Phi(x^{\ast}(p^{\dagger})),$ (10)
where $J(x,p)$ is the game Jacobian defined as $J_{i}(x,p)=\nabla_{{x}_{i}}\ell_{i}(x)+p_{i}$ for every $i\in\mathcal{I}$. Next, under Assumption 2.1, the variational inequality characterization of Nash equilibrium [FP07], we know that $x^{\ast}(p^{\dagger})$ is a Nash equilibrium if and only if
$\langle J(x^{\ast}(p^{\dagger}),p^{\dagger}),{x}-x^{\ast}(p^{\dagger})\rangle\geqslant 0,\quad\forall\ {x}\in X.$ (11)
From (10) and (11) observe that
$\langle\nabla\Phi(x^{\ast}(p^{\dagger})),{x}-x^{\ast}(p^{\dagger})\rangle\geqslant 0,\quad\forall\ {x}\in X.$ (12)
Further, from the first order conditions of optimality for social cost function we know that $x^{\dagger}$ is socially optimal if any only if it satisfies
$\langle\nabla\Phi(x^{\dagger}),x-x^{\dagger}\rangle\geqslant 0,\quad\forall\ x\in X.$ (13)
Comparing (12) with (13), we note that $x^{\ast}(p^{\dagger})$ is the minimizer of social cost function $\Phi$. This implies $x^{\ast}(p^{\dagger})={x}^{\dagger}$ as ${x}^{\dagger}$ is the unique minimizer of social cost function $\Phi$ by Assumption 2.1.
Similarly for non-atomic game $\tilde{G}$, we show that the incentive $\tilde{p}^{\dagger}$ aligns the Nash equilibrium with social optimality (i.e. for any $\tilde{p}^{\dagger}\in\tilde{{P}}^{\dagger}$, ${\tilde{x}^{\ast}}(\tilde{p}^{\dagger})={\tilde{x}}^{\dagger}$). Fix $\tilde{p}^{\dagger}\in\tilde{{P}}^{\dagger}$. For every $j\in\tilde{S}_{i}$ and $i\in\tilde{\mathcal{I}}$, it holds that $\tilde{p}^{j\dagger}_{i}=\tilde{e}_{i}^{j}({\tilde{x}^{\ast}}(\tilde{p}^{\dagger}))$. This implies that
$\tilde{c}_{i}^{j}({\tilde{x}^{\ast}}(\tilde{p}^{\dagger}),\tilde{p}^{\dagger})=\nabla_{\tilde{x}_{i}^{j}}\tilde{\Phi}({\tilde{x}^{\ast}}(\tilde{p}^{\dagger})).$ (14)
Under Assumption 2.2, the variational inequality characterization of Nash equilibrium for non-atomic games [San10], we know that ${\tilde{x}^{\ast}}(\tilde{p}^{\dagger})$ is a Nash equilibrium if and only if
$\langle\tilde{c}({\tilde{x}^{\ast}}(\tilde{p}^{\dagger}),\tilde{p}^{\dagger}),\tilde{x}-{\tilde{x}^{\ast}}(\tilde{p}^{\dagger})\rangle\geqslant 0\quad\forall\ \tilde{x}\in\tilde{X}.$ (15)
From (14) and (15), we observe that
$\langle\nabla\tilde{\Phi}({\tilde{x}^{\ast}}(\tilde{p}^{\dagger})),\tilde{x}-{\tilde{x}^{\ast}}(\tilde{p}^{\dagger})\rangle\geqslant 0\quad\forall\ \tilde{x}\in\tilde{X}.$ (16)
Recall from the first order conditions of optimality for social cost function, we know that $\tilde{x}^{\dagger}$ is socially optimal if any only if it satisfies
$\langle\nabla\tilde{\Phi}(\tilde{x}^{\dagger}),\tilde{x}-\tilde{x}^{\dagger}\rangle\geqslant 0,\quad\forall\ \tilde{x}\in\tilde{X}.$ (17)
Comparing (16) with (17), we note that ${\tilde{x}^{\ast}}(\tilde{p}^{\dagger})$ is the minimizer of social cost function $\tilde{\Phi}$. This implies ${\tilde{x}^{\ast}}(\tilde{p}^{\dagger})={\tilde{x}}^{\dagger}$ as ${\tilde{x}}^{\dagger}$ is the unique minimizer of social cost function $\tilde{\Phi}$ by Assumption 2.2.
Next, we show that the set $P^{\dagger}$ is singleton. We prove this via contradiction. Suppose $P^{\dagger}$ contains two element $p^{\dagger}_{1},p^{\dagger}_{2}$, which are shown to align the Nash equilibrium with social optimality. Therefore, it holds that $x^{\dagger}=x^{\ast}(p^{\dagger}_{1})=x^{\ast}(p^{\dagger}_{2})$. Moreover, due to (LABEL:eq:_PDagger) we know that $p^{\dagger}_{1}=e(x^{\ast}(p^{\dagger}_{1}))$ and $p^{\dagger}_{2}=e(x^{\ast}(p^{\dagger}_{2}))$. Consequently, we have $p^{\dagger}_{1}=e(x^{\dagger})=p^{\dagger}_{2}$. The proof of uniqueness of $\tilde{P}^{\dagger}$ also follows analogously. ∎
Remark 3.2.
Proposition 3.1 demonstrates that the externality-based incentive updates ($p$-update) and ($\tilde{p}$-update) have two advantages. Firstly, these updates guarantee that any fixed point must achieve social optimality. Secondly, the result indicates that the optimal incentive mechanism takes a simple format: the incentive mechanism is linear in players’ strategies in atomic games and is a constant vector in non-atomic games.
3.2 Convergence to optimal incentive mechanism
In this subsection, we provide sufficient conditions for convergence of strategy and incentive updates ($x$-update)-($p$-update) and ($\tilde{x}$-update)-($\tilde{p}$-update). In order to study the convergence of these updates we make the following assumption:
Assumption 3.1.
The step sizes in ($x$-update)-($p$-update) and ($\tilde{x}$-update)-($\tilde{p}$-update) satisfy the following condition
  • (i)
    $\sum_{k=1}^{\infty}\gamma_{k}=\sum_{k=1}^{\infty}\beta_{k}=+\infty$, $\sum_{k=1}^{\infty}\gamma_{k}^{2}+\beta_{k}^{2}<+\infty$.
  • (ii)
    $\lim_{k\to\infty}\frac{\beta_{k}}{\gamma_{k}}=0$.
In Assumption 3.1, (i) is a standard assumption on step-sizes that allows us to analyze the convergence of the discrete-time learning updates via a continuous-time differential equation [Bor97]. Additionally, (ii) assumes that the incentive update occurs at a slower timescale compared to the update of strategies.
In view of Assumption 3.1-(ii), it is necessary that the strategy updates converge if the incentives are held fixed. This condition holds true for a variety of updates commonly studied in learning in games literature. For instance, the equilibrium update (cf. (4)) converges to Nash equilibrium in one-step. Similarly, the best response update (cf. (5)) converges in zero sum games [Har98], potential games [SMK18], and dominance solvable games [Nac90]. The gradient-based update (cf. (6)) converge in concave games [Ros65]. Therefore, in this work we do not focus on studying the convergence of strategy updates with fixed incentive mechanisms and we pick any off-the-shelf convergent strategy update. Our goal is to characterize conditions under which the coupled strategy and incentive updates ($x$-update)-($p$-update) and ($\tilde{x}$-update)-($\tilde{p}$-update) converge. Therefore, we impose the following assumption which is sufficient to ensure convergence of strategy updates ($x$-update) and ($\tilde{x}$-update) with fixed incentive mechanisms. Particularly, we assume that if the incentive mechanism is fixed, then the corresponding continuous-time dynamical system converges to the Nash equilibrium associated with that incentive mechanism.
Assumption 3.2.
For any incentive mechanism $p,\tilde{p}$, consider the following continuous time dynamical systems associated with ($x$-update) and ($\tilde{x}$-update), respectively,
$\dot{x}(t)$ $=f(x(t),p)-x(t),$ ($x$-dynamics)
$\dot{\tilde{x}}(t)$ $=\tilde{f}(\tilde{x}(t),\tilde{p})-\tilde{x}(t).$ ($\tilde{x}$-dynamics)
The equilibrium $x^{\ast}(p)$ (resp. $\tilde{x}^{\ast}(\tilde{p})$) is the global asymptotically stable fixed point of the dynamics ($x$-dynamics) (resp. ($\tilde{x}$-dynamics)).
Using the standard approximation arguments from [Bor09] it is evident that Assumption 3.1 and 3.2 are sufficient to ensure convergence of strategy updates ($x$-update) and ($\tilde{x}$-update) with fixed incentive mechanisms.
Against the preceding backdrop, in the following result, we characterize the convergence conditions for the coupled strategy and incentive updates.
Proposition 3.3.
Consider an atomic game $G$. Suppose that Assumptions 2.1, 3.1, 3.2 and at least one of the following conditions holds:
  • (C1)
    $\frac{\partial e_{i}(x^{\ast}(p))}{\partial p_{j}}>0$ for all $p\in\mathbb{R}^{n}$ and all $i\neq j$. Additionally, at least one of the following conditions hold:
    • (i)
      If $e_{i}(x^{\ast}(0))\geqslant 0$ for every $i\in\mathcal{I}$ then there exists $p\in\mathbb{R}_{+}^{|\mathcal{I}|}$ such that $e_{i}(x^{\ast}(p))-p_{i}\leqslant 0$ for every $i\in\mathcal{I}$. Moreover, $x_{1}\in X,p_{1}\in\mathbb{R}^{|\mathcal{I}|}_{+}$.
    • (ii)
      If $e_{i}(x^{\ast}(0))\leqslant 0$ for every $i\in\mathcal{I}$ then there exists $p\in\mathbb{R}_{-}^{|\mathcal{I}|}$ such that $e_{i}(x^{\ast}(p))-p_{i}\geqslant 0$ for every $i\in\mathcal{I}$. Moreover, $x_{1}\in X,p_{1}\in\mathbb{R}^{|\mathcal{I}|}_{-}$.
  • (C2)
    There exists a set $\textsf{dom}(V)\subset\mathbb{R}^{|\mathcal{I}|}$ and a continuously differentiable function $V:\textsf{dom}(V)\rightarrow\mathbb{R}_{+}$ such that $V(p^{\dagger})=0$ and $V(p)>0$ for all $p\neq p^{\dagger}$. Moreover:
    $\nabla V(p)^{\top}\left(e(x^{\ast}(p))-p\right)<0\quad\forall\ p\neq p^{\dagger}.$
Additionally, let $\sup_{k\in\mathbb{N}}(\|x_{k}\|+\|p_{k}\|)<+\infty$. Then there exist positive scalars $\bar{r},\bar{\alpha},\bar{\beta},\bar{\gamma}$ such that the discrete-time updates ($x$-update)-($p$-update) in $G$ when initialized at $p_{0}\in\mathcal{B}_{\bar{r}}(p^{\dagger})$, $x_{0}\in\mathcal{B}_{\bar{r}}(x^{\ast}(p_{0}))$, with step-sizes such that
$\sup_{k\in\mathbb{N}}\beta_{k}/\gamma_{k}\leqslant\bar{\alpha},\quad\sup_{k\in\mathbb{N}}\beta_{k}\leqslant\bar{\beta},\quad\sup_{k\in\mathbb{N}}\gamma_{k}\leqslant\bar{\gamma},$ (18)
satisfy [^2] The results in this paper hold even if the updates ($x$-update)-($p$-update) and ($\tilde{x}$-update)-($\tilde{p}$-update) are perturbed with martingale difference noise [BP18].
$\lim_{k\to\infty}(x_{k},p_{k})=({x}^{\dagger},p^{\dagger}).$ (19)
The same argument holds for non-atomic game $\tilde{G}$ under Assumptions 2.2, 3.1, and 3.2.
Proof.
Under Assumption 3.1, the two-timescale approximation theory [Bor97, BP18] suggests that the strategy update ($x$-update) is a fast transient while the incentive update ($p$-update) is a slow component. This separation in timescales allow us to study the convergence of these updates in two stages. First, we study the convergence of fast strategy updates, for every fixed value of incentive. Second, we study the convergence of slow incentive updates, assuming that the fast strategy updates have converged to the equilibrium.
Formally, to study the convergence of fast strategy updates, we re-write ($x$-update)-($p$-update) as follows
$x_{k+1}$ $=x_{k}+\gamma_{k}\left(f(x_{k},p_{k})-x_{k}\right)$ (20)
$p_{k+1}$ $=p_{k}+\gamma_{k}\frac{\beta_{k}}{\gamma_{k}}\left(e(x_{k})-p_{k}\right).$
Since $\sup_{k\in\mathbb{N}}(\|x_{k}\|+\|p_{k}\|)<+\infty$ and $\lim_{k\rightarrow\infty}\beta_{k}/\gamma_{k}=0$ (cf. Assumption 3.1), the term $\frac{\beta_{k}}{\gamma_{k}}\left(e(x_{k})-p_{k}\right)$ in (20) goes to zero as $k\rightarrow\infty$. Consequently, leveraging the standard approximation arguments [Bor09, Lemma 1, Section 2.2], we conclude that the asymptotic behavior of the updates in (20) is same as that of the following dynamical system
$\dot{\textbf{x}}(t)=f(\textbf{x}(t),\textbf{p}(t))-\textbf{x}(t),\quad\dot{\textbf{p}}(t)=0.$
Using Assumption 3.2, we conclude that
$\lim_{k\rightarrow\infty}({x}_{k},p_{k})\rightarrow\{(x^{\ast}(p),p):p\in\mathbb{R}^{|\mathcal{I}|}\}.$ (21)
Next, to study the convergence of slow incentive updates, we re-write ($p$-update) as follows
$p_{k+1}=p_{k}+\beta_{k}\left(e(x^{\ast}(p_{k}))-p_{k}\right)+\beta_{k}\left(e(x_{k})-e(x^{\ast}(p_{k}))\right).$ (22)
We will show that $(p_{k})_{k\in\mathbb{N}}$ will asymptotically follow the trajectories of the following continuous-time dynamical system:
$\dot{\textbf{p}}(t)$ $=e(x^{\ast}(\textbf{p}(t)))-\textbf{p}(t).$ (23)
Note that $p^{\dagger}$ is the fixed point of the trajectories of dynamical system (23) (cf. Proposition 3.1). Conditions (C1) and (C2) in Proposition 3.3 are based on results from non-linear dynamical systems which ensure convergence of (23). In particular, (C1)-(i) (resp. (C1)-(ii)) builds on the cooperative dynamical systems theory [Hir85], which ensures that $\mathbb{R}_{+}^{|\mathcal{I}|}$ (resp. $\mathbb{R}_{-}^{|\mathcal{I}|}$) is positively invariant for (23) and $p^{\dagger}\in\mathbb{R}_{+}^{|\mathcal{I}|}$ (resp. $p^{\dagger}\in\mathbb{R}_{-}^{|\mathcal{I}|}$) is asymptotically stable. On the other hand, condition (C2) ensures the existence of a Lyapunov function that is strictly positive everywhere except at $p^{\dagger}$ and decreases along the flow of (23) ([Sas13]). Either condition guarantees the convergence of the trajectories of (23) to $p^{\dagger}$.
Let $D^{\dagger}$ denote the domain of attraction of $p^{\dagger}$ for the dynamical system (23). From converse Lyapunov theorem [Sch64], we know that there exists a continuously differentiable function $\bar{V}:D^{\dagger}\rightarrow\mathbb{R}_{+}$ such that $\bar{V}(p^{\dagger})=0$, $\bar{V}(p)>0$ for all $p\in D^{\dagger}\backslash\{p^{\dagger}\}$ and $\bar{V}(p)\rightarrow\infty$ as $p\rightarrow\textsf{boundary}(D^{\dagger})$. For any $r>0$, define $\bar{V}_{r}=\{p\in\textsf{dom}(\bar{V}):\bar{V}(p)\leqslant r\}$ to be a sub-level set of $\bar{V}$. There exists $0<\bar{r}^{\prime}<\bar{r}$ such that $\bar{V}_{\bar{r}^{\prime}}\subsetneq\mathcal{B}_{\bar{r}^{\prime}}(p^{\dagger})\subsetneq\mathcal{B}_{\bar{r}}(p^{\dagger})\subsetneq\bar{V}_{\bar{r}}$. Additionally, define $\ell_{0}=0,\ell_{k}=\sum_{i=1}^{k}\beta_{i}$ and $L_{k}=\ell_{n(k)}$ where $n(0)=0$, and
$n(k)=\min\left\{m\geqslant n(k-1):\sum_{j=n(k-1)+1}^{m}\beta_{j}\geqslant T\right\}~{}\forall k\in\mathbb{N}.$ (24)
Here, $T$ is a positive integer to be described shortly. Furthermore, define $\bar{\textbf{p}}^{(k)}:\mathbb{R}_{+}\rightarrow\mathbb{R}^{|\mathcal{I}|}$ to be a solution of (23) on $[L_{k},\infty)$ such that $\bar{\textbf{p}}^{(k)}(L_{k})=p_{L_{k}}$.
To ensure that $\bar{\textbf{p}}^{(k)}(L_{k})\in\textsf{dom}(\bar{V})$ for $k>0$, we show that for an appropriate choice of $T$ in (24), $p_{L_{k}}\in\textsf{int}(D^{\dagger})$ for every $k\in\mathbb{N}$. From [BP18, Theorem IV.1], we know that there exists $K>0$ such that for all $k\in\mathbb{N},$
$\|p_{k}-\bar{\textbf{p}}^{(0)}(\ell_{k})\|$
$\leqslant K\left(\sup_{k}\beta_{k}+\sup_{k}\gamma_{k}+\sup_{k}\frac{\beta_{k}}{\gamma_{k}}+\sup_{k}\frac{\beta_{k}}{\gamma_{k}}\|x_{0}-x^{\ast}(p_{0})\|\right)$
$=K\left(\bar{\alpha}+\bar{\beta}+\bar{\gamma}+\bar{\alpha}\bar{r}\right)=:\kappa.$
Consequently, using triangle inequality, it holds that
$\|p_{k}-p^{\dagger}\|\leqslant\kappa+\|\bar{\textbf{p}}^{(0)}(\ell_{k})-p^{\dagger}\|.$ (25)
Since $\bar{V}$ is a Lyapunov function of (23) and $\bar{\textbf{p}}^{(0)}(0)=p_{0}\in\mathcal{B}_{\bar{r}}(p^{\dagger})\subsetneq\bar{V}_{\bar{r}}$, there exists $\bar{k}\in\mathbb{N}$ such that for all $k\geqslant\bar{k}$, $\bar{\textbf{p}}^{(0)}(\ell_{k})\in\bar{V}_{\bar{r}^{\prime}}\subsetneq\mathcal{B}_{\bar{r}^{\prime}}(p^{\dagger})$. If we choose $\kappa<\bar{r}-\bar{r}^{\prime}$ then, from (25), it holds that for all $k\geqslant\bar{k}$, $p_{k}\in\mathcal{B}_{\bar{r}}(p^{\dagger})$. Therefore, if we choose $T\geqslant\bar{k}$ in (24), it holds that
$p_{L_{k}}\in\textsf{dom}(\bar{V}),\quad\forall\ k\in\mathbb{N}.$ (26)
Define $\hat{p}:\mathbb{R}_{+}\rightarrow\mathbb{R}$ such that, for every $k\in\mathbb{N}$, $\hat{p}(\ell_{k})=p_{k}$ with linear interpolation on $[\ell_{k},\ell_{k+1}]$. Using the standard approximation arguments from [Bor09, Chapter 6], it holds that333 For any $T\geqslant\bar{k}$ and $\delta>0$, there exists $k(\delta)$ such that $\hat{p}(\ell_{k(\delta)}+\cdot)$ form a “$(T,\delta)$” perturbation (cf. [Bor97]) of (23).
$\sup_{t\in[L_{k},L_{k+1}]}\|\hat{p}(t)-\bar{\textbf{p}}^{(k)}(t)\|$
$\leqslant\mathcal{O}\left(\sum_{m\geqslant L_{k}}\beta_{m}^{2}+\sup_{m\geqslant L_{k}}\|x_{m}-x^{\ast}(p_{m})\|\right).$ (27)
Using (21) and Assumption 3.1, we conclude that RHS in the above equation goes to zero as $k\rightarrow\infty$. Finally, using (26), (3.2) and [Bor97, Lemma 2.1] we conclude that $p_{k}\rightarrow p^{\dagger}$ as $k\rightarrow\infty$. ∎
4 Applications
In this section, we study the applicability of the proposed dynamic incentive design approach in two practically relevant class of games: atomic aggregative games, and non-atomic routing games.
4.1 Atomic Aggregative Games
Here, we study quadratic network aggregative games [BK07, BGR16, SCM22a, AJ13]. Consider a game $G$ comprised of a finite set of players $\mathcal{I}$. The strategy set of every player is the entire real line $\mathbb{R}$. Given the joint strategy profile $x=(x_{i})_{i\in\mathcal{I}}$, the cost of each player $i\in\mathcal{I}$ is
$\ell_{i}(x)=\frac{1}{2}q_{i}x_{i}^{2}+\alpha x_{i}(Ax)_{i},$ (28)
where $A\in\mathbb{R}^{|\mathcal{I}|\times|\mathcal{I}|}$ is the network matrix, where $A_{ij}$ represents the impact of player $j$’s strategy on the cost of player $i$, and $\alpha>0$ is a constant that characterizes the impact of the aggregate strategy on the individual cost of players. Moreover, $q_{i}>0$ characterizes the impact of each player’s own strategy on their cost function. Without loss of generality, we consider $A_{ii}=0$ for all $i\in\mathcal{I}$. For notational brevity, we define $Q=\textsf{diag}((q_{i})_{i\in\mathcal{I}})\in\mathbb{R}^{|\mathcal{I}|\times|\mathcal{I}|}$.
A social planner designs incentives by setting a payment $p_{i}x_{i}$ for player $i$ for choosing strategy $x_{i}$. Thus, given a incentive vector $p=(p_{i})_{i\in\mathcal{I}}$, the total cost of player $i$ is given by $c_{i}(x,p)=\ell_{i}(x)+p_{i}x_{i}$. The social planner’s objective is to ensure that the players reach a strategy $x^{\dagger}$, which minimizes the following social cost function:
$\Phi(x)=\sum_{i=1}^{n}h_{i}(x_{i}),$ (29)
Assumption 4.1.
We assume that for every $i\in\mathcal{I}$, $h_{i}(\cdot)$ is a strictly convex function with gradient being Lipschitz continuous. Furthermore, we assume existence of $y^{\dagger}\in\mathbb{R}^{|\mathcal{I}|}$ such that $\nabla h_{i}(y_{i}^{\dagger})=0$ for every $i\in\mathcal{I}$.
Such separable social costs have been considered in the literature on control of agents in aggregative games [Gra17] as well as systemic risk analysis in financial networks [AOTS15]. One example of $\Phi(x)$ that satisfies this property include $\Phi(x)=\frac{1}{2}\sum_{i\in\mathcal{I}}(x_{i}-\xi_{i})^{2}$, which computes the difference between the players’ strategies and some desired strategy $\xi\in\mathbb{R}^{|\mathcal{I}|}$. Another example is $\Phi(x)=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}x_{i}$ representing the aggregate level of all agents’ activities in the network [AOTS15].
Proposition 4.1.
Suppose that Assumption 4.1 holds and $M:=Q+\alpha A$ is invertible. Then Nash equilibrium ${x^{\ast}}(p)$ exists and is unique for every incentive vector $p\in\mathbb{R}^{|\mathcal{I}|}$. Furthermore, the set $P^{\dagger}$ is a singleton set.
Proof.
First, we show that if $M$ is invertible, then ${x^{\ast}}(p)=-M^{-1}p$ for any $p\in\mathbb{R}^{|\mathcal{I}|}$. To see this, observe that the cost function $c_{i}(x_{i},x_{-i},p)$ is strongly convex in $x_{i}$ and the strategy space $X_{i}$ is unconstrained. Therefore, it must hold that $\nabla_{x_{i}}c_{i}(x^{\ast}(p),p)=0$, for every $i\in\mathcal{I}$. Consequently, using (28), it must hold that
$q_{i}x_{i}^{\ast}(p)+\alpha(Ax^{\ast}(p))_{i}+p_{i}=0,\quad\forall\ i\in\mathcal{I}.$ (30)
Stacking (30) in a vector form yields $Mx^{\ast}(p)=-p$.
Next, we show that $P^{\dagger}$ is a singleton set. From (7a), the externality in this setup is expressed as:
$e_{i}(x)=\nabla h_{i}(x_{i})-q_{i}x_{i}-\alpha\sum_{j=1}^{n}A_{ij}x_{j},~{}\forall i\in\mathcal{I},~{}\forall x.$ (31)
Combining (30) and (31), we obtain
$e_{i}(x^{\ast}(p))=\nabla h_{i}(x_{i}^{\ast}(p))+p_{i}.$ (32)
Consequently, using (LABEL:eq:_PDagger), we obtain
$P^{\dagger}=\{p^{\dagger}\in\mathbb{R}^{|\mathcal{I}|}:\nabla h_{i}(x^{\ast}_{i}(p^{\dagger}))=0,~{}\ \forall\ i\in\mathcal{I}\}.$ (33)
Since $h_{i}$ is strictly convex, from Assumption 4.1, there exists a unique $y^{\dagger}$ such that $\nabla h_{i}(y_{i}^{\dagger})=0$, for every $i\in I$. Therefore, for every $p^{\dagger}\in P^{\dagger}$ it should hold that $x^{\ast}(p^{\dagger})=y^{\dagger}$. Using the fact that $x^{\ast}(p)=-M^{-1}p$, we obtain the unique $p^{\dagger}=-My^{\dagger}$. ∎
Next, we provide sufficient conditions to ensure the convergence of ($x$-update)-($p$-update).
Proposition 4.2.
Consider the aggregative game $G$ with invertible $M$. Suppose that Assumptions 3.1, 3.2, and 4.1 are satisfied, and at least one of the following conditions hold:
  • (A1)
    The matrix $M$ has non-negative entries, $M^{-1}$ has strictly-negative off-diagonal entries. Furthermore, there exists [^4] Since $h_{i}$ is strictly convex then $\nabla h_{i}(\cdot)$ is strictly increasing and can have only one zero-crossing, which ensures that $y^{\dagger}$ is unique. some $y^{\dagger}\in\mathbb{R}^{|\mathcal{I}|}$ such that $\nabla h_{i}(y^{\dagger}_{i})=0$ for every $i\in\mathcal{I}$, and all entries of $y^{\dagger}$ have the same sign;
  • (A2)
    $M$ is a symmetric positive definite matrix.
Additionally, $\sup_{k\in\mathbb{N}}(\|x_{k}\|+\|p_{k}\|)<+\infty$. Then there exist positive scalars $\bar{r},\bar{\alpha},\bar{\beta},\bar{\gamma}$ such that for every $p_{0}\in\mathcal{B}_{\bar{r}}(p^{\dagger})$, $x_{0}\in\mathcal{B}_{\bar{r}}(x^{\ast}(p_{0}))$, with step-sizes that follow (18) must satisfy $\lim_{k\to\infty}(x_{k},p_{k})=({x}^{\dagger},p^{\dagger})$.
Proof.
The proof follows by verifying that the conditions $(A1)$ and $(A2)$ ensure that the conditions $(C1)$ and $(C2)$ in Proposition 3.3 are satisfied, respectively.
To begin, we show that condition $(A1)$ ensures that condition $(C1)$ is satisfied in Proposition 3.3. Recall, from (32), that $e_{i}(x^{\ast}(p))=\nabla h_{i}(x_{i}^{\ast}(p))+p_{i}$. First, we show that $i,j\in\mathcal{I}$ such that $i\neq j,$ $\frac{\partial e_{i}(x^{\ast}(p))}{\partial p_{j}}>0$. Indeed,
$\frac{\partial e_{i}(x^{\ast}(p))}{\partial p_{j}}$ $=\nabla^{2}h_{i}(x_{i}^{\ast}(p))\frac{\partial x_{i}^{\ast}(p)}{\partial p_{j}}$
$=\nabla^{2}h_{i}(x_{i}^{\ast}(p))(-M^{-1})_{ij}>0,$
where the inequality holds because $h_{i}$ is strictly convex and $(M^{-1})_{ij}<0$.
Next, we show that if $e_{i}(x^{\ast}(0))\geqslant 0$ for every $i\in\mathcal{I}$ then there exists $p\in\mathbb{R}_{+}^{|\mathcal{I}|}$ such that $e_{i}(x^{\ast}(p))-p_{i}\leqslant 0$ for every $i\in\mathcal{I}$ (cf. (C1)-(i) in Proposition 3.3). Using (32), it is sufficient to show that
[…]
$\nabla h_{i}(x^{\ast}(\bar{p}))\leqslant 0\quad\forall\ i\in\mathcal{I}.$
By $(A1)$, we know that $\nabla h_{i}(y^{\dagger}_{i})=0$ and $y^{\dagger}_{i}y^{\dagger}_{j}\geqslant 0$ for every $i,j\in\mathcal{I}$. We claim that $y^{\dagger}_{i}\nabla h_{i}(0)\leqslant 0$ for every $i\in\mathcal{I}$. Indeed, for every $i\in\mathcal{I}$,
$0\leqslant(\nabla h_{i}(y^{\dagger}_{i})-\nabla h_{i}(x^{\ast}_{i}(0)))(y^{\dagger}_{i}-x_{i}^{\ast}(0))=-y^{\dagger}_{i}\nabla h_{i}(0),$
where the first inequality is due to the strict convexity of $h_{i}$. Suppose $\nabla h_{i}(0)\geqslant 0$ , for every $i\in\mathcal{I}$, then $-y^{\dagger}\in\mathbb{R}_{+}^{|\mathcal{I}|}$. Next, define $\bar{p}=-(1+\epsilon)My^{\dagger}$, for some $\epsilon>0$, then $\bar{p}\in\mathbb{R}_{+}^{|\mathcal{I}|}$ as $y^{\dagger}\in\mathbb{R}_{+}^{|\mathcal{I}|}$ and $M$ has all non-negative entries (cf. $(A1)$). This ensures that $x^{\ast}(\bar{p})=(1+\epsilon)y^{\dagger}$. We claim that $\nabla h_{i}(x^{\ast}_{i}(\bar{p}))<0$ for every $i\in\mathcal{I}$. Indeed, for every $i\in\mathcal{I}$,
00 0 $<(\nabla h_{i}(x_{i}^{\ast}(\bar{p}))-\nabla h_{i}(y^{\dagger}_{i}))(x_{i}^{\ast}(\bar{p})-y^{\dagger}_{i})$
$=\nabla h_{i}(x_{i}^{\ast}(\bar{p}))\epsilon y^{\dagger}_{i}.$
The claim follows because $y^{\dagger}_{i}\leqslant 0$ and $\epsilon>0$. This shows that the implication in (34) holds. Similarly, it can be show that if $e_{i}(x^{\ast}(0))\leqslant 0$ for every $i\in\mathcal{I}$ then there exists $p\in\mathbb{R}_{-}^{|\mathcal{I}|}$ such that $e_{i}(x^{\ast}(p))-p_{i}\geqslant 0$ for every $i\in\mathcal{I}$ (cf. (C1)-(ii) in Proposition 3.3).
Next, we verify that condition $(A2)$ ensures that condition $(C2)$ in Proposition 3.3 is satisfied. Consider the function $V(p)=(p-p^{\dagger})^{\top}M^{-\top}(p-p^{\dagger})$. Note that $V(p^{\dagger})=0$ and $V(p)>0$ for all $p\neq p^{\dagger}$. Next, we show that $\nabla V(p)^{\top}(e(x^{\ast}(p))-p)\leqslant 0.$ Indeed,
$\nabla V(p)^{\top}(e(x^{\ast}(p))-p)=2(p-p^{\dagger})^{\top}M^{-\top}\nabla h(x^{\ast}(p))$
$=-2(x^{\ast}(p)-x^{\ast}(p^{\dagger}))\nabla h(x^{\ast}(p))$
$=-2(x^{\ast}(p)-x^{\ast}(p^{\dagger}))\left(\nabla h(x^{\ast}(p))-\nabla h(x^{\ast}(p^{\dagger}))\right)$
$=-2(x^{\ast}(p)-x^{\ast}(p^{\dagger}))\left(\nabla h(x^{\ast}(p))-\nabla h(x^{\ast}(p^{\dagger}))\right)$
$\leqslant 0$
where the last equality holds due to the strict convexity of $h_{i}$ for every $i\in\mathcal{I}$. ∎
Remark 4.3.
There exists instances of aggregative game such that only one of the conditions $(A1)$ or $(A2)$ is satisfied in Proposition 4.2. For instance consider the matrices
$M_{1}=\begin{bmatrix}1&0.1\\ 1&1\end{bmatrix},M_{2}=\begin{bmatrix}1&-0.1\\ -0.1&1\end{bmatrix}.$
Note that $M_{1}$ satisfies $(A1)$ but not $(A2)$ and the matrix $M_{2}$ satisfies $(A2)$ but not $(A1)$.
4.2 Non-atomic Traffic Routing on General Networks
Consider a routing game $\tilde{G}$ characterizing the interaction of strategic travelers over a graph $\tilde{\mathcal{G}}=(\tilde{\mathcal{E}},\tilde{\mathcal{N}})$ where $\tilde{\mathcal{N}}$ is the set of nodes and $\tilde{\mathcal{E}}$ denotes the set of edges. Let $\tilde{\mathcal{I}}$ be the set of origin-destination (o-d) pairs. Each o-d pair $i\in\tilde{\mathcal{I}}$ is connected by a set of routes [^5] A route is a set of contiguous edges. represented by $\tilde{R}_{i}$. Let $\tilde{R}=\cup_{i\in\tilde{\mathcal{I}}}\tilde{R}_{i}$ denote the set of all routes on the network.
An infinitesimal traveler on the network is associated with an o-d pair and chooses a route to commute between the o-d pair. Let the total population of travelers associated with any o-d pair $i\in\tilde{\mathcal{I}}$ be denote by $\tilde{M}_{i}$. Let $\tilde{x}_{i}^{j}$ be the amount of travelers taking route $j\in\tilde{R}_{i}$ to commute between o-d pair $i\in\tilde{\mathcal{I}}$ and $\tilde{x}=(\tilde{x}_{i}^{j})_{j\in\tilde{R}_{i},i\in\tilde{\mathcal{I}}}$ is a vector which contains, as its entries, the route flow of all population on different routes. Naturally, it holds that $\sum_{j\in\tilde{R}_{i}}\tilde{x}_{i}^{j}=\tilde{M}_{i}$ for every $i\in\tilde{\mathcal{I}}$. Any route flow $\tilde{x}$ induces a flow on the edges of the network denoted by $\tilde{w}$ such that $\tilde{w}_{a}=\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}\tilde{x}_{i}^{j}\mathbbm{1}(a\in j)$ for every $a\in\tilde{\mathcal{E}}$. We denote the set of feasible route flows by $\tilde{X}$ and the set of feasible edge flows by $\tilde{W}=\{(\tilde{w}_{a})_{a\in\tilde{\mathcal{E}}}:\exists\tilde{x}\in\tilde{X},\tilde{w}_{a}=\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}\tilde{x}_{i}^{j}\}$.
For any o-d pair $i\in\tilde{\mathcal{I}}$ and route flow $\tilde{x}\in\mathbb{R}^{|\tilde{R}|}$, the cost experienced by travelers using route $j\in\tilde{R}_{i}$ is ~$\tilde{\ell}_{i}^{j}(\tilde{x})=\sum_{a\in\tilde{\mathcal{E}}}l_{a}(\tilde{w}_{a})\mathbbm{1}(a\in j)$, where $l_{a}(\cdot)$ is the edge latency function that depends on the edge flows. For every edge $a\in\tilde{\mathcal{E}}$, we assume that the edge latency function $l_{a}(\cdot)$ is convex and strictly increasing. This property of edge latency function captures the congestion effect on the transportation network [BMW56, Rou10].
A social planner designs incentives by setting tolls on the edges of the network in the form of edge tolls [^6] If we directly use the setup of non-atomic games presented in Section 2.1.2 we would require the social planner to use route-based tolls rather than edge-based tolls. Our approach of using edge-based tolls is rooted in practical consideration with implementation of tolls., denoted by $\tilde{p}=(\tilde{p}_{a})_{a\in\tilde{\mathcal{E}}}$. Every edge toll vector induces a unique route toll vector $\tilde{P}$. That is, for any o-d pair $i\in\tilde{\mathcal{I}}$ the toll on route $j\in\tilde{R}_{i}$ is
$\tilde{P}_{i}^{j}=\sum_{a\in\tilde{\mathcal{E}}:a\in j}\tilde{p}_{a}.$ (35)
Consequently, the total cost experienced by travelers on o-d pair $i\in\tilde{\mathcal{I}}$ who choose route $j\in\tilde{R}_{i}$ is $\tilde{c}_{i}^{j}(\tilde{x},\tilde{P})=\tilde{\ell}_{i}^{j}(\tilde{x})+\tilde{P}_{i}^{j}$. Let $\tilde{x}^{\ast}(\tilde{P})$ denote a Nash equilibrium corresponding to route tolls $\tilde{P}$. Owing to (35), with slight abuse of notation, we shall frequently use $\tilde{x}^{\ast}(\tilde{P})$ and $\tilde{x}^{\ast}(\tilde{p})$ interchangeably. Typically, the equilibrium route flows can be non-unique but the corresponding edge flows $\tilde{w}^{\ast}(\tilde{p})$ are unique. Furthermore, the function $\tilde{p}\mapsto\tilde{w}^{\ast}(\tilde{p})$ is a continuous function [YH05].
The social planner’s objective is to design tolls that ensure that the resulting equilibrium minimizes the overall travel time incurred by travelers on the network, characterized as the minimizer of
$\tilde{\Phi}(\tilde{x})=\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}\tilde{x}_{i}^{j}\tilde{\ell}_{i}^{j}(\tilde{x}).$ (36)
Note that the optimal route flow can be non-unique but the optimal edge flow, denoted by $w^{\dagger}$, is unique [YH05].
Using the description of cost of travelers, the externality, based on (7b), caused by travelers from o-d pair $i\in\tilde{\mathcal{I}}$ using the route $j\in\tilde{R}_{i}$ is
$\tilde{e}_{i}^{j}(\tilde{x})=\sum_{i^{\prime}\in\tilde{\mathcal{I}}}\sum_{j^{\prime}\in\tilde{R}_{i}}\tilde{x}_{i^{\prime}}^{j^{\prime}}\frac{\partial\tilde{\ell}_{i^{\prime}}^{j^{\prime}}(\tilde{x})}{\partial\tilde{x}_{i}^{j}}\stackrel{{\scriptstyle(a)}}{{=}}\sum_{a\in\tilde{\mathcal{E}}:a\in j}\nabla l_{a}(\tilde{w}_{a})\tilde{w}_{a},$ (37)
where $(a)$ is due to Lemma B.2 in Appendix B.
Define
$\tilde{\textbf{P}}^{\dagger}=\{(\tilde{p}^{\dagger}_{a})_{a\in\tilde{\mathcal{E}}}:\tilde{p}^{\dagger}_{a}=\tilde{w}^{\ast}_{a}(\tilde{p}^{\dagger})\nabla l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}^{\dagger})),\forall\ a\in\tilde{\mathcal{E}}\big{\}}.$
Proposition 4.4.
The set $\tilde{\textbf{P}}^{\dagger}$ is non-empty singleton set. The unique $p^{\dagger}\in\tilde{\textbf{P}}^{\dagger}$ is socially optimal, i.e. $\tilde{w}(p^{\dagger})=w^{\dagger}$.
Proof.
First, we show that $\tilde{\textbf{P}}^{\dagger}$ is non-empty. This can be shown analogous to the proof of existence in Proposition 3.1 by using Schauder fixed point theorem and continuity of function $\tilde{w}^{\ast}(\cdot)$. We omit the details of this proof for the sake of brevity.
Next, we show that any $p^{\dagger}\in\tilde{\textbf{P}}^{\dagger}$ aligns the Nash equilibrium with social optimality, i.e. $\tilde{w}(p^{\dagger})=w^{\dagger}$. For any $p^{\dagger}\in\tilde{\textbf{P}}^{\dagger}$, we have $\tilde{p}^{\dagger}_{a}=\tilde{w}^{\ast}_{a}(\tilde{p}^{\dagger})\nabla l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}^{\dagger}))$ for every $a\in\tilde{\mathcal{E}}$. This implies,
$\frac{\partial}{\partial\tilde{w}_{a}}(\tilde{w}_{a}(\tilde{p}^{\dagger})l_{a}(\tilde{w}_{a}(\tilde{p}^{\dagger})))=l_{a}(\tilde{w}_{a}(\tilde{p}^{\dagger}))+\tilde{p}^{\dagger}_{a},\quad\forall\ a\in\tilde{\mathcal{E}}.$ (38)
Note that for any arbitrary edge toll $\tilde{p}\in\mathbb{R}^{|\tilde{\mathcal{E}}|}$, $\tilde{w}^{\ast}(\tilde{p})$ is the unique solution to the following strictly convex optimization problem [San10].
[…]
Therefore, $\tilde{w}^{\ast}(\tilde{p})$ is a Nash equilibrium if and only if
$\sum_{a\in\tilde{\mathcal{E}}}\left(l_{a}(\tilde{w}_{a}(\tilde{p})+\tilde{p}_{a})(\tilde{w}_{a}-\tilde{w}_{a}(\tilde{p}))\right)\geqslant 0\quad\forall\ \tilde{w}\in\tilde{W}.$ (40)
Combining (38) and (40), we conclude that
$\sum_{a\in\tilde{\mathcal{E}}}\frac{\partial}{\partial\tilde{w}_{a}}(\tilde{w}_{a}(\tilde{p}^{\dagger})l_{a}(\tilde{w}_{a}(\tilde{p}^{\dagger})))(\tilde{w}_{a}-\tilde{w}_{a}^{\ast}(\tilde{p}^{\dagger}))\geqslant 0\quad\forall\ \tilde{w}\in\tilde{W}.$ (41)
Further, from the first order conditions of optimality for social cost function, we know that $\tilde{x}^{\dagger}$ is socially optimal if any only if
$\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}\frac{\partial\Phi(\tilde{x}^{\dagger})}{\partial\tilde{x}_{i}^{j}}(\tilde{x}_{i}^{j}-\tilde{x}_{i}^{\dagger j})\geqslant 0.$ (42)
Using Lemma B.3, we can equivalently write (42) in terms of edge flows as follows
$\sum_{a\in\tilde{\mathcal{E}}}\frac{\partial}{\partial\tilde{w}_{a}}(\tilde{w}_{a}^{\dagger}l_{a}(\tilde{w}_{a}^{\dagger}))(\tilde{w}_{a}-\tilde{w}_{a}^{\dagger})\geqslant 0\quad\forall\ \tilde{w}\in\tilde{W},$ (43)
where $w^{\dagger}$ is the edge flow corresponding to the route flow $x^{\dagger}$. Comparing (41) with (43), we note that $\tilde{w}^{\ast}(p^{\dagger})$ is the minimizer of social cost function $\tilde{\Phi}$. This implies $\tilde{w}^{\ast}(\tilde{p}^{\dagger})=\tilde{w}^{\dagger}$.
The proof that $\tilde{\textbf{P}}^{\dagger}$ is singleton follows by contradiction which is analogous to that in Proposition 3.1. We omit the details for the sake of brevity.
Next, we show that the updates ($\tilde{x}$-update)-($\tilde{p}$-update) converge to the fixed points discussed in previous section.
Proposition 4.5.
Suppose that for every $i\in\tilde{I},j\in\tilde{R}_{i}$
$\tilde{c}_{i}^{j}(\tilde{x}^{\ast}(\tilde{p}^{\dagger}))\leqslant\tilde{c}_{i}^{j^{\prime}}(\tilde{x}^{\ast}(\tilde{p}^{\dagger}))\ \forall\ j^{\prime}\in\tilde{R}_{i}\implies\tilde{x}_{i,j}^{\ast}(\tilde{p}^{\dagger})>0.$ (44)
Additionally, $\sup_{k\in\mathbb{N}}(\|\tilde{x}_{k}\|+\|\tilde{p}_{k}\|)<+\infty$. Then there exist positive scalars $\bar{r},\bar{\alpha},\bar{\beta},\bar{\gamma}$ such that for every $\tilde{p}_{0}\in\mathcal{B}_{\bar{r}}(\tilde{p}^{\dagger})$, $\tilde{x}_{0}\in\mathcal{B}_{\bar{r}}(\tilde{x}^{\ast}(\tilde{p}_{0}))$, with step-sizes that follow (18) must satisfy $\lim_{k\to\infty}(\tilde{x}_{k},\tilde{p}_{k})=(\tilde{x}^{\dagger},\tilde{p}^{\dagger})$.
Proof.
The proof follows by verifying the second convergence condition in Proposition 3.3. Particularly, we show that $V(\tilde{p})=(\tilde{p}-\tilde{p}^{\dagger})^{\top}\Delta(\tilde{p}-\tilde{p}^{\dagger})$ serves as candidate Lyapunov function which satisfies the conditions stated in Proposition 3.3-(C2) if we set $\Delta\in\mathbb{R}^{|\tilde{\mathcal{E}}|\times|\tilde{\mathcal{E}}|}$ to be a diagonal matrix such that
$\Delta_{a,a}=(\nabla l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}^{\dagger}))+\tilde{w}^{\ast}_{a}(\tilde{p}^{\dagger})\nabla^{2}l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}^{\dagger})))^{-1}.$
Due to the strict monotonicity and convexity of $l_{a}(\cdot)$, it follows that $\Delta_{a,a}>0$ for every $a\in\tilde{\mathcal{E}}$. Towards this goal, we show that there exists a positive scalar $r$ such that for any $\tilde{p}\in\mathcal{B}_{r}(\tilde{p}^{\dagger})$, it holds that
$\sum_{a\in\tilde{\mathcal{E}}}\nabla_{\tilde{p}_{a}}V(\tilde{p})^{\top}\left(\tilde{w}^{\ast}_{a}(\tilde{p})\nabla l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}))-\tilde{p}_{a}\right)<-2V(\tilde{p}).$ (45)
Indeed, we note that
$\sum_{a\in\tilde{\mathcal{E}}}\nabla_{\tilde{p}_{a}}V(\tilde{p})\left(\tilde{w}^{\ast}_{a}(\tilde{p})\nabla l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}))-\tilde{p}_{a}\right)$
$=2\sum_{a\in\tilde{\mathcal{E}}}\Delta_{a,a}(\tilde{p}_{a}-\tilde{p}^{\dagger}_{a})\left(\tilde{w}^{\ast}_{a}(\tilde{p})\nabla l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}))-\tilde{p}_{a}\right)$
$=2\sum_{a\in\tilde{\mathcal{E}}}\Delta_{a,a}(\tilde{p}_{a}-\tilde{p}^{\dagger}_{a})\left(\tilde{w}^{\ast}_{a}(\tilde{p})\nabla l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}))-\tilde{p}^{\dagger}_{a}+\tilde{p}^{\dagger}_{a}-\tilde{p}_{a}\right)$
$=-2V(\tilde{p})+2\sum_{a\in\tilde{\mathcal{E}}}\Delta_{a,a}(\tilde{p}_{a}-\tilde{p}^{\dagger}_{a})\left(\phi_{a}(\tilde{p})-\phi_{a}(\tilde{p}^{\dagger})\right)$
where for every $a\in\tilde{\mathcal{E}}$, $\phi_{a}(\tilde{p}):=\tilde{w}^{\ast}_{a}(\tilde{p})\nabla l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}))$. Thus it suffices to show that
$\sum_{a\in\tilde{\mathcal{E}}}\Delta_{a,a}(\tilde{p}_{a}-\tilde{p}^{\dagger}_{a})\left(\phi_{a}(\tilde{p})-\phi_{a}(\tilde{p}^{\dagger})\right)\leqslant 0.$ (46)
Condition (44) ensures that the function $\tilde{\phi}$ is differentiable in a neighborhood of $\tilde{p}^{\dagger}$ (cf. [YH05, Chapter 4]). From Lemma B.1, we know that if $z^{\top}\Delta D\phi(\tilde{p}^{\dagger})z\leqslant 0$ for all $z$ then there exists $r>0$ such that (46) holds for all $\tilde{p}\in\mathcal{B}_{r}(\tilde{p}^{\dagger})$. Indeed, by design of $\Delta$ it holds that $\Delta D\phi(\tilde{p}^{\dagger})=D\tilde{w}(\tilde{p}^{\dagger})$. Finally, using Lemma B.1 and Lemma B.4, we note that $z^{\top}D\tilde{w}(\tilde{p}^{\dagger})z\leqslant 0$. This concludes the proof. ∎
5 Concluding Remarks
We propose an adaptive incentive mechanism that $(i)$ updates based on the feedback of each agent’s externality, $(ii)$ is agnostic to the learning rule employed by agents, and $(iii)$ updates at a slower timescale compared to the agents’ learning dynamics, leading to a two-timescale coupled dynamical system. We provide sufficient conditions that ensure the convergence of this dynamical system to the optimal incentive mechanism, guaranteeing that the Nash equilibrium of the corresponding game among agents achieves social optimality. We demonstrate the effectiveness of our proposed incentive mechanism in two practically relevant games: atomic networked quadratic aggregative games and non-atomic network routing games, by verifying the proposed convergence conditions.
An important direction for future research is to verify the convergence conditions in a broader class of games. Notably, leveraging the externality-based dynamic incentive design update proposed in an earlier version of this paper (cf. [MKWS21]), [LMZ23] studied the convergence of such updates in power systems.
Appendix A Counter-example.
In this section, we present an instance of a non-atomic game where the standard gradient-based incentive design approach would fail. Specifically, we will show that the social cost, as a function of the incentive (evaluated at Nash equilibrium), is non-convex and has a zero gradient, which would result in the gradient-based approach not converging.
Consider a non-atomic routing game, comprising of two nodes and two edges connecting them. This network is used by one unit of travelers traveling from the source node $\mathcal{S}$ to the destination node $\mathcal{D}$. The latency function of two edges are denoted in Figure 1.
$\mathcal{S}$$\mathcal{D}$$c_{1}(\tilde{x},\tilde{p})=\tilde{x}^{1}+\tilde{p}^{1}$$c_{2}(\tilde{x},\tilde{p})=\tilde{x}^{2}+\tilde{p}^{2}$
Figure 1: Two-link routing game.
In this game, the strategy set is $\tilde{X}=\{\tilde{x}\in\mathbb{R}^{2}:\tilde{x}_{1}+\tilde{x}_{2}=1\}$. The equilibrium congestion levels on the two edges is obtained by computing the minimizer of the following function [YH05]:
$T(\tilde{x},\tilde{p})=\frac{1}{2}\tilde{x}_{1}^{2}+\frac{1}{2}\tilde{x}_{2}^{2}+\tilde{p}_{1}\tilde{x}_{1}+\tilde{p}_{2}\tilde{x}_{2}.$
Thus, for any toll vector $\tilde{p}$ the Nash equilibrium is ${\tilde{x}^{\ast}}(\tilde{p})=\arg\min_{\tilde{x}\in\tilde{X}}T(\tilde{x},\tilde{p})$ is such that ${\tilde{x}^{\ast}_{1}}(\tilde{p})=\mathcal{P}_{[0,1]}\left(\frac{\tilde{p}_{2}-\tilde{p}_{1}+1}{2}\right),{\tilde{x}^{\ast}_{2}}(\tilde{p})=\mathcal{P}_{[0,1]}\left(\frac{\tilde{p}_{1}-\tilde{p}_{2}+1}{2}\right),$ where for any scalar $x\in\mathbb{R}$, $\mathcal{P}_{[0,1]}(x)$ denotes its projection on the line segment $[0,1]$.
The social cost function (cf. (36)) is given by $\tilde{\Phi}(\tilde{x})=\tilde{x}_{1}\tilde{\ell}_{1}(\tilde{x})+\tilde{x}_{2}\tilde{\ell}_{2}(\tilde{x})=\tilde{x}_{1}^{2}+\tilde{x}_{2}^{2}$. Thus,
$\tilde{\Phi}({\tilde{x}^{\ast}}(\tilde{p}))$ =$=\begin{cases}\frac{(\tilde{p}_{1}-\tilde{p}_{2})^{2}+1}{2}&\text{ if }|\tilde{p}_{1}-\tilde{p}_{2}|\leqslant 1,\\ 1&\text{ otherwise }\end{cases}$
However, this function is not convex, for example, on the set $|\tilde{p}_{1}-\tilde{p}_{2}|<1.5$. Therefore, gradient based updates initialized in the region $\{(\tilde{p}_{1},\tilde{p}_{2})\in\mathbb{R}^{2}:1.1\leqslant|\tilde{p}_{1}-\tilde{p}_{2}|\leqslant 1.5\}$ will not converge to the global optimizer as the gradient is zero in this region. Thus, the approach from [LNL+23, LLY+21] does not apply here.
Appendix B Auxiliary Results
Lemma B.1 ([FP07]).
For any fixed $p^{\prime}$ and continuously differentiable function ϕ:~~:italic-$\phi:\mathbb{R}^{\tilde{\mathcal{E}}}\rightarrow\mathbb{R}^{\tilde{\mathcal{E}}}$ the condition
$\langle\phi(p)-\phi(p^{\prime}),p-p^{\prime}\rangle\leqslant 0\quad\forall\ p\in\mathcal{B}_{r}(p^{\prime})$
for some $r>0$, holds if and only if $z^{\top}D\phi(p^{\prime})z\leqslant 0$ for all $z\in\mathbb{R}^{|\tilde{\mathcal{E}}|}$.
Lemma B.2.
The externality $\tilde{e}_{i}^{j}$ in (37) can be alternatively written, in terms of edge flows, as $\tilde{e}_{i}^{j}(\tilde{x})=\sum_{a\in j}\tilde{w}_{a}\nabla l_{a}(\tilde{w}_{a}).$
Proof.
Expanding the externality expression from (37) we observe that
$\tilde{e}_{i}^{j}(\tilde{x})$ $=\sum_{i^{\prime}\in\tilde{\mathcal{I}}}\sum_{j^{\prime}\in\tilde{R}_{i}}\tilde{x}_{i^{\prime}}^{j^{\prime}}\frac{\partial\tilde{\ell}_{i^{\prime}}^{j^{\prime}}(\tilde{x})}{\partial\tilde{x}_{i}^{j}}$
=$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{i^{\prime}\in\tilde{\mathcal{I}}}\sum_{j^{\prime}\in\tilde{R}_{i}}\tilde{x}_{i^{\prime}}^{j^{\prime}}\sum_{a\in\tilde{\mathcal{E}}}\mathbbm{1}(a\in j^{\prime})\frac{\partial l_{a}(\tilde{w}_{a})}{\partial\tilde{x}_{i}^{j}}$
=$\stackrel{{\scriptstyle(b)}}{{=}}\sum_{i^{\prime}\in\tilde{\mathcal{I}}}\sum_{j^{\prime}\in\tilde{R}_{i}}\tilde{x}_{i^{\prime}}^{j^{\prime}}\sum_{a\in\tilde{\mathcal{E}}}\mathbbm{1}(a\in j^{\prime})\nabla l_{a}(\tilde{w}_{a})\frac{\partial\tilde{w}_{a}}{\partial\tilde{x}_{i}^{j}}$
=$\stackrel{{\scriptstyle(c)}}{{=}}\sum_{i^{\prime}\in\tilde{\mathcal{I}}}\sum_{j^{\prime}\in\tilde{R}_{i}}\tilde{x}_{i^{\prime}}^{j^{\prime}}\sum_{a\in\tilde{\mathcal{E}}}\mathbbm{1}(a\in j^{\prime})\nabla l_{a}(\tilde{w}_{a})\mathbbm{1}(a\in j)$
=$\stackrel{{\scriptstyle(d)}}{{=}}\sum_{a\in\tilde{\mathcal{E}}}\nabla l_{a}(\tilde{w}_{a})\mathbbm{1}(a\in j)\sum_{i^{\prime}\in\tilde{\mathcal{I}}}\sum_{j^{\prime}\in\tilde{R}_{i}}\tilde{x}_{i^{\prime}}^{j^{\prime}}\mathbbm{1}(a\in j^{\prime})$
=$\stackrel{{\scriptstyle(e)}}{{=}}\sum_{a\in j}\nabla l_{a}(\tilde{w}_{a})\tilde{w}_{a}$
where $(a)$ follows by expanding out the expression of route costs in terms of edge costs, $(b)$ follows by the chain rule, $(c)$ follows due to the definition of edge flows, $(d)$ follows by change of order of summations and finally $(e)$ follows by the definition of edge flows. This completes the proof. ∎
Lemma B.3.
$x^{\dagger}$ that satisfies (42) if and only if (43).
Proof.
First, we show that
$\Phi(\tilde{x})=\sum_{a\in\tilde{\mathcal{E}}}\tilde{w}_{a}l_{a}(\tilde{w}_{a}).$ (47)
Indeed,
[…]
$=\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}\tilde{x}_{i}^{j}\sum_{a\in\tilde{\mathcal{E}}}\mathbbm{1}(a\in j)l_{a}(\tilde{w}_{a})$
$=\sum_{a\in\tilde{\mathcal{E}}}l_{a}(\tilde{w}_{a})\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}\tilde{x}_{i}^{j}\mathbbm{1}(a\in j)=\sum_{a\in\tilde{\mathcal{E}}}l_{a}(\tilde{w}_{a})\tilde{w}_{a}.$
Next, observe that
$\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}\frac{\partial\Phi(\tilde{x}^{\dagger})}{\partial\tilde{x}_{i}^{j}}(\tilde{x}_{i}^{j}-\tilde{x}_{i}^{\dagger j})$
$=\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}\sum_{a\in\tilde{\mathcal{E}}}\frac{\partial}{\partial\tilde{x}_{i}^{j}}\left(\tilde{w}_{a}l_{a}(\tilde{w}_{a})\right)(\tilde{x}_{i}^{j}-\tilde{x}_{i}^{\dagger j})$
$=\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}\sum_{a\in\tilde{\mathcal{E}}}\frac{\partial}{\partial\tilde{w}_{a}}\left(\tilde{w}_{a}l_{a}(\tilde{w}_{a})\right)\mathbbm{1}(a\in j)(\tilde{x}_{i}^{j}-\tilde{x}_{i}^{\dagger j})$
$=\sum_{a\in\tilde{\mathcal{E}}}\frac{\partial}{\partial\tilde{w}_{a}}\left(\tilde{w}_{a}l_{a}(\tilde{w}_{a})\right)\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}\mathbbm{1}(a\in j)(\tilde{x}_{i}^{j}-\tilde{x}_{i}^{\dagger j})$
$=\sum_{a\in\tilde{\mathcal{E}}}\frac{\partial}{\partial\tilde{w}_{a}}\left(\tilde{w}_{a}l_{a}(\tilde{w}_{a})\right)(\tilde{w}_{a}-\tilde{w}_{a}^{\dagger}).$
This concludes the proof. ∎
Lemma B.4.
For any $\tilde{p},\tilde{p}^{\prime}\in\mathbb{R}^{|\tilde{\mathcal{E}}|}$ the following holds
$\sum_{a\in\tilde{\mathcal{E}}}(\tilde{p}_{a}-\tilde{p}^{\prime}_{a})(\tilde{w}^{\ast}_{a}(\tilde{p}_{a})-\tilde{w}^{\ast}_{a}(\tilde{p}^{\prime}_{a}))\leqslant 0.$ (48)
Proof.
To prove this result, we first show that
$\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}(\tilde{P}_{i}^{j}-\tilde{P}_{i}^{{}^{\prime}j})(\tilde{x}_{i}^{\ast j}(\tilde{P})-\tilde{x}_{i}^{\ast j}(\tilde{P}^{\prime}))\leqslant 0,$ (49)
where $\tilde{P}$ and $\tilde{P}^{\prime}$ are the route tolls associated with edge tolls $\tilde{p}$ and $\tilde{p}^{\prime}$ respectively through (35).
Let the feasible set of the optimization problem (39) be denoted by $\mathcal{F}$. Using the first-order conditions of optimality for the strictly convex optimization problem (39) we obtain the following variational inequality:
$\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}$ $\left(\tilde{c}_{i}^{j}(\tilde{x}^{\ast}(\tilde{P}),\tilde{P})\right)\cdot\left(\tilde{y}_{i}^{j}-\tilde{x}_{i}^{\ast j}(\tilde{P})\right)\geqslant 0,\forall\quad\tilde{y}\in\mathcal{F},$ (50)
where $\tilde{P}$ is the route toll associated with edge toll $\tilde{p}$. Rewriting (50) for edge tolls $\tilde{p}^{\prime}$ we obtain
$\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}$ $\left(\tilde{c}_{i}^{j}(\tilde{x}^{\ast}(\tilde{P}^{\prime}),\tilde{P}^{\prime})\right)\cdot\left(\tilde{y}_{i}^{{}^{\prime}j}-\tilde{x}_{i}^{\ast j}(\tilde{P}^{\prime})\right)\geqslant 0,\forall\quad\tilde{y}^{\prime}\in\mathcal{F},$ (51)
where $\tilde{P}^{\prime}$ is the route toll associated with edge toll $\tilde{p}^{\prime}$.
We are now ready to show the condition (49). Note that
$\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}(\tilde{P}_{i}^{j}-\tilde{P}_{i}^{{}^{\prime}j})(\tilde{x}_{i}^{\ast j}(\tilde{P})-\tilde{x}_{i}^{\ast j}(\tilde{P}^{\prime}))$
$\stackrel{{\scriptstyle(a)}}{{\leqslant}}\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}(\tilde{\ell}_{i}^{j}(\tilde{x}^{\ast}(\tilde{P}^{\prime}))-\tilde{\ell}_{i}^{j}(\tilde{x}^{\ast}(\tilde{P})))(\tilde{x}_{i}^{\ast j}(\tilde{P})-\tilde{x}_{i}^{\ast j}(\tilde{P}^{\prime}))$
=$\stackrel{{\scriptstyle(b)}}{{=}}\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}(\tilde{x}_{i}^{\ast j}(\tilde{P})-\tilde{x}_{i}^{\ast j}(\tilde{P}^{\prime}))\cdot$
$\hskip 56.9055pt\cdot\sum_{a\in\tilde{\mathcal{E}}}(l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}^{\prime}))-l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p})))\mathbbm{1}(a\in j)$
=$\stackrel{{\scriptstyle(c)}}{{=}}\sum_{a\in\tilde{\mathcal{E}}}(l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}^{\prime}))-l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p})))\cdot$
$\hskip 56.9055pt\cdot\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}(\tilde{x}_{i}^{\ast j}(\tilde{P})-\tilde{x}_{i}^{\ast j}(\tilde{P}^{\prime}))\mathbbm{1}(a\in j)$
=$\stackrel{{\scriptstyle(d)}}{{=}}\sum_{a\in\tilde{\mathcal{E}}}(l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}^{\prime}))-l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p})))(\tilde{w}^{\ast}_{a}(\tilde{p})-\tilde{w}^{\ast}_{a}(\tilde{p}^{\prime}))$
$\stackrel{{\scriptstyle(e)}}{{\leqslant}}0$ (52)
where we obtain $(a)$ by adding (50), evaluated at $\tilde{y}=\tilde{x}^{\ast}(\tilde{P}^{\prime})$, and (51), evaluated at $\tilde{y}^{\prime}=\tilde{x}^{\ast}(\tilde{P})$, $(b)$ holds by definition of route loss function, $(c)$ holds by interchange of summation, $(d)$ holds by definition of edge flows and $(e)$ holds due to monotonicity of edge latency functions. This proves (49).
Now we are ready to prove (48). Note that
$\sum_{a\in\tilde{\mathcal{E}}}(\tilde{p}_{a}-\tilde{p}_{a^{\prime}})(\tilde{w}^{\ast}_{a}(\tilde{p})-\tilde{w}^{\ast}_{a}(\tilde{p}^{\prime}))$
=$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{a\in\tilde{\mathcal{E}}}(\tilde{p}_{a}-\tilde{p}_{a^{\prime}})\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}(\tilde{x}_{i}^{\ast j}(\tilde{P})-\tilde{x}_{i}^{\ast j}(\tilde{P}^{\prime}))\mathbbm{1}(a\in j)$
=$\stackrel{{\scriptstyle(b)}}{{=}}\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}(\tilde{x}_{i}^{\ast j}(\tilde{P})-\tilde{x}_{i}^{\ast j}(\tilde{P}^{\prime}))\sum_{a\in\tilde{\mathcal{E}}}(\tilde{p}_{a}-\tilde{p}_{a^{\prime}})\mathbbm{1}(a\in j)$
=$\stackrel{{\scriptstyle(c)}}{{=}}\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}(\tilde{x}_{i}^{\ast j}(\tilde{P})-\tilde{x}_{i}^{\ast j}(\tilde{P}^{\prime}))(\tilde{P}_{i}^{j}-\tilde{P}_{i}^{{}^{\prime}j})$
$\stackrel{{\scriptstyle(d)}}{{\leqslant}}0,$
where $(a)$ holds due to the definition of edge flows, $(b)$ holds due to interchange of summation, $(c)$ holds due to definition of route tolls, $(d)$ holds due to (49). This concludes the proof. ∎